# Robust and Smooth Estimation of the Extreme Tail Index via Weighted Minimum Density Power Divergence Saida Mancer, Abdelhakim Necir Djamel

### v1

@@ 1. Introduction
The widely adopted approach for estimating the parameters of an extreme value distribution in extreme value analysis is the maximum likelihood estimation (MLE) method. While MLEs enjoy desirable asymptotic properties, they can be highly sensitive to outliers or deviations from the assumed extreme value models Brazauskas and Serfling (2000). Robust statistics offer a valuable alternative for mitigating the influence of outliers, and consequently, reducing deviations from the underlying parametric models. Incorporating robust statistical methods into extreme value theory has been shown to enhance both the accuracy and reliability of parameter estimates. (Dell’Aquila and Embrechts, 2006). Juárèz and Schucany (2004) appear to be the first to employ the Minimum Density Power Divergence (MDPD) introduced by Basu et al. (1998) for the robust estimation of parameters in extreme value distributions. Since then, this divergence measure has become one of the most widely used tools for robust parameter estimation in the context of extreme value theory. Several authors, including Kim and Lee (2008), Dierchx et al. (2013), Goegebeur et al. (2014), Dierckx et al. (2021) have employed the MDPD framework to estimate the tail index and quantiles of Pareto-type distributions.
@@ 3. Weighted MDPD estimation of the tail index
By substituting with and with in the previous equation, we en up with the WMDPD estimator of tail index as solution of the equation

### v2

@@ 1. Introduction
The widely adopted approach for estimating the parameters of an extreme value distribution in extreme value analysis is the maximum likelihood estimation (MLE) method. While MLEs enjoy desirable asymptotic properties, they can be highly sensitive to outliers or deviations from the assumed extreme value models Brazauskas and Serfling (2000). Robust statistics offer a valuable alternative for mitigating the influence of outliers, and consequently, reducing deviations from the underlying parametric models. Incorporating robust statistical methods into extreme value theory has been shown to enhance both the accuracy and reliability of parameter estimates (Dell’Aquila and Embrechts, 2006). Juárèz and Schucany (2004) appear to be the first to employ the Minimum Density Power Divergence (MDPD) introduced by Basu et al. (1998) for the robust estimation of parameters in extreme value distributions. Since then, this divergence measure has become one of the most widely used tools for robust parameter estimation in the context of extreme value theory. Several authors, including Kim and Lee (2008), Dierchx et al. (2013), Goegebeur et al. (2014), Dierckx et al. (2021) have employed the MDPD framework to estimate the tail index and quantiles of Pareto-type distributions.
@@ 3. Weighted MDPD estimation of the tail index
By substituting with and with in the previous equation, we end up with the WMDPD estimator of tail index as solution of the equation